OpenStudy (anonymous):

2n^2-30n^2+100n=0 Solve pls help

OpenStudy (anonymous):

did you by chance type that wrong ?

OpenStudy (anonymous):

Well looking at the problem, you can pull out an n n(2n-30n+100)=0 Simplify the inside n(-28n+100)=0 and set each to 0 n=0 (-28n+100)=0 So the answers are 0 and whatever the thing below it is.